OpenStudy (anonymous):

f(x)+x^4[f(x)]^3=1028 and f(2)=4, find f'(2)

OpenStudy (bahrom7893):

differentiate implicitly

OpenStudy (bahrom7893):

f'(x) + 4x^3*(f(x))^3 + 3(f(x))^2*f '(x)+x^4 = 0

OpenStudy (bahrom7893):

Plug in x = 2: f '(2) + 4*8 * 4^3 + 3 * 4^2 * f'(2) + 2^4 = 0

OpenStudy (bahrom7893):

f'(2) + 2048 + 48 f'(2) + 16 = 0

OpenStudy (bahrom7893):

49f'(2) = -2048-16 49 f'(2) = -2064 f'(2) = -42.2.. hmm i might've made a careless mistake somewhere, jsut check

OpenStudy (anonymous):

thanks very much for the help bahrom, can you help me with this last one use implicit differentiation to find equation of tangent line to -1x^2+3xy+4y^3=-8 at the point (4,2)

OpenStudy (anonymous):

in y=mx + b form

OpenStudy (bahrom7893):

i wonder what ash has been typing lol... let's let him finish

OpenStudy (anonymous):

he left wow

OpenStudy (anonymous):

my software rejected the -42.2 so i guess i would have to check

OpenStudy (anonymous):

can you help me with the other question i posted pls

OpenStudy (bahrom7893):

yea sure.. for the first one i might have made some careless mistake, just check the product rule and multiplication and values

OpenStudy (ash2326):

We have \[f(x)+x^4[f(x)]^3=1028\] Let's differentiate above equation with respect to x \[f'(x)+4x^3{f(x)}^3+3x^4 f(x)^2 f'(x)=0\] let's substitute x=2 we get \[f'(2)+4\times 2^3 f(2)^3 +3*2^4f(2)^2f'(2)=0\] f(2)=4 so we get \[f'(2)+4\times8 \times 64+3\times 16\times 16 f'(2)=0\] we have \[f'(2)+2048+4096 f'(2)=0\] \[4097 f'(2)=-2048\] \[f'(2)=-2048/4097\]

OpenStudy (bahrom7893):

3(f(x))^2*f '(x)*x^4 it was supposed to be times..

OpenStudy (bahrom7893):

-1x^2+3xy+4y^3=-8 at the point (4,2) -2x + 3xy'+3y + 12y^2 y' = 0

OpenStudy (bahrom7893):

3xy' + 12y^2y' = 2x - 3y y' (3x+12y^2) = 2x - 3y y' = (2x-3y)/(3x+12y^2) y' = (8 - 6)/(12+48) = 2/60 = 1/30

OpenStudy (bahrom7893):

y = (1/30)x + b

OpenStudy (bahrom7893):

now we just gotta figure out when what y is when x=0

OpenStudy (bahrom7893):

for that use: (y-y1)/(x-x1) = 1/30 (4-y1)/(2-0) = 1/30 4-y1 = 1/15 4 - 1/15 = y1 y1 = 59/15 y = (1/30)x+(59/15) Again i hate the numbers, check for careless mistakes

OpenStudy (anonymous):

neither work, but i will go back to check

OpenStudy (bahrom7893):

ughhh hold on.. ill write this out.. this is annoying

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

For the first one try -2048/769

OpenStudy (anonymous):

yea that works, thanks

OpenStudy (bahrom7893):

okay.. lol i thought i was going mad..

OpenStudy (anonymous):

lol

OpenStudy (bahrom7893):

all right the slope of the 2nd one is correct

OpenStudy (anonymous):

ok. my software is rejecting. perhaps, its just how am I typing it in. I may need to figure out how to correctly input it.

OpenStudy (bahrom7893):

is it called webworks? lol

OpenStudy (bahrom7893):

no i might be doing something wrong

OpenStudy (anonymous):

yea, does your school use it as well?

OpenStudy (bahrom7893):

well i was lucky enough to skip it.. but i used to solve my friends' webworks for them lol

OpenStudy (anonymous):

lucky you. i hate this thing. sometimes correctly entering your solution is more difficult than solving the actual problem

OpenStudy (bahrom7893):

i know.. i once decided to round 2.999 up to 3... god the answer it finally accepted was 2.999

OpenStudy (anonymous):

lol i have encounter that issue alot of times

OpenStudy (bahrom7893):

found my mistake

OpenStudy (anonymous):

nice. where did you go wrong?

OpenStudy (bahrom7893):

4-y1 = 1/15 4 - 1/15 = y1 y1 = 59/15 y = (1/30)x+(59/15) i used x instead of y and y instead of x.

OpenStudy (anonymous):

oh ok

OpenStudy (bahrom7893):

y = (1/30)x+(28/15) <- your answer

OpenStudy (bahrom7893):

Please click on Good Answer if I helped. You can also hover over my name and Become a fan ;)!